Saturday 06/28/14

10:45 am – 5:30 pm, 9:00 pm – 11:50 pm

Small tasks

  • restart data copying after computer freeze last night (GPU computation abort failed)
  • restart sticky polymer simulations

STORM imaging

  • imaging sample C10 (blue 33 kb vg locus).
  • Conv images nice and bright
  • images look pretty good.
  • STORM imaging sample C11 (black 10 kb)

Meeting with Ajaz

  • follow up to ANTP PH cluster analysis: BX-C and control region
  • Ajaz found new potential polymer modeling collaborator (see email)
  • will get manuscript to XZ

Chromatin Data analysis

  • Analyzed D01 data (ChromatinCropper)
  • Analyzed D02 data
  • what happened to D03 data?!
  • Analyzed D04 data
  • Analyzed D05 data
  • Analyzed D06 data (way too large, not sure what’s up with this locus!)
  • Analyzed D07 data
  • black trends looking decent

Cell culture

  • passaged cells yesterday
  • start new cultures from frozen stock today
  • density looking pretty good, should be able to plate a batch of these cells tomorrow and passage the rest.

Chromatin experiment planning

  • ordered toe-hold displacement probes to try to remove un-bleached, un-imaged secondaries for sequential hybe approach.
This entry was posted in Summaries and tagged , . Bookmark the permalink.